You are planning to paint a house. A) How many square feet will you need to paint? (there are four windows, each 6 feet by 3 feet; two windows, each 2 feet by 30 feet; and two doors, each 80/36, that do not require paint.).b() the paint that you have chosen is available in gallon can only. Each can covers 900 square feet. If you want to use two coats of paint how many cans will you need for the project? C) if the paint you have chosen sells for $24.31 per gallon, what will it cost to paint the house?



the house has two heights listed which are 40 and 30


the width is 50ft and the length is 70 ft

Answer by addingup(3677)

Are you painting the ceiling? If the problem and/or the teacher didn't say, assume you are only painting the walls.
It's a rectangular room, 50 x 70, so you have
2*50*? = (the ? stands for the height of the walls, you need to decide)
2*70*? =
-------------
Total = ?
- * - * - * - * - * - *
For your reference, 30 or 40 for the height of the walls seems awfully high, you'd better double check.
- * - * - * - * - * - *
Now calculate all openings (windows and doors) that you don't have to paint:
(4*6*3)+(2*2*30)+(2*80/36) = Subtract this from your total for the walls
- * - * - * - * - * - *
NOTE!!: What is 80/36 for the doors?? Is it 80 x 36? Those would be HUGE doors. Are you mixing feet and inches?
- * - * - * - * - * - *
The rest of the problem you can solve yourself easily with a calculator, just get this part correctly first.